A 75 kg. block of wood rest on the top of the smooth plane surface, whose length is 4 meters and whose altitude is 1.0 meter. How long will it take for th block to slide down to the bottom of the plane when released?



1
Expert's answer
2022-04-19T12:05:53-0400

The speed of the block at the beginning of motion is "v_i =0m\/s". Ssnce the surface is smooth, the speed at the end of the motion "v_f" can be find from the energy conservation law:


"mgh=\dfrac{mv_f^2}{2}"

where "m= 75kg, h = 1.0m, g= 9.8m\/s^2". Thus, obtain:


"v_f = \sqrt{2gh}"

Next, using the kinematic formula, find the time of motion "t" :


"d = \dfrac{v_i+v_f}{2}\cdot t"

where "d = 4m". Thus, obtain:


"t = \dfrac{2d}{v_f} = \dfrac{2d}{\sqrt{2gh}} = d\sqrt{\dfrac{2}{gh}}\\
t = 4\cdot \sqrt{\dfrac{2}{9.8\cdot1}} \approx 2s"

Answer. 2s.